OpenStudy (anonymous):

Find the sum of the sequence 39+40+41+42+....+138

OpenStudy (klimenkov):

Firstly, tell me how many terms does this sequence have?

OpenStudy (anonymous):

First term a = 39 and last term l = 138 Use: \[t_n = a + (n-1)d\] find n from it and then put it into: \[\huge Sum = \frac{n}{2}(a + l)\]

OpenStudy (anonymous):

There, \(t_n = 138\)

OpenStudy (anonymous):

138 = 39 + (n-1)(1) 138 = 38 + n n = 100 Just plug in the formula above: \[Sum = \frac{100}{2}(39 + 138) = 50 \times 177 = 8850\]